# A Flow-Based Model for Conditional and Probabilistic Electricity Consumption Profile Generation and Prediction

### v1

@@ Abstract
Residential Load Profile (RLP) generation and prediction are critical for the operation and planning of distribution networks, particularly as diverse low-carbon technologies are increasingly integrated. This paper introduces a novel flow-based generative model, termed Full Convolutional Profile Flow (FCPFlow), which is uniquely designed for both conditional and unconditional RLP generation, and for probabilistic load forecasting. By introducing two new layers–the invertible linear layer and the invertible normalization layer–the proposed FCPFlow architecture shows three main advantages compared to traditional statistical and contemporary deep generative models: 1) it is well-suited for RLP generation under continuous conditions, such as varying weather and annual electricity consumption, 2) it shows superior scalability in different datasets compared to traditional statistical, and 3) it also demonstrates better modeling capabilities in capturing the complex correlation of RLPs compared with deep generative models.
@@
load profile, flow-based model, generative model, probabilistic load prediction.
@@ I Introduction
RLPs have wide applications in areas such as energy supply and demand management [1], modern system planning[2], and risk analysis [3]. The validity of these studies depends largely on the quality of RLPs used. However, access to RLP data is limited due to privacy [4]. RLP generation can provide effective solutions to these problems. On the one hand, distribution system operators (DSOs) rely on RLPs for refining planning decisions. Historical data inaccessibility or limitations can hinder this process, yet the generation of RLPs provides system planners with an alternative to executing more informed planning [5]. For instance, in [6], generated RLPs were used to understand consumption patterns and optimize the system planning. On the other hand, generated RLPs can function as augmented data to support high-level tasks. For example, generated RLPs or PV profiles are used to support the training of models for load prediction [7], non-intrusive load monitoring algorithms [8], and reinforcement learning [9].
@@
Traditional RLP modeling primarily employs Gaussian Mixture Models (GMMs) [10, 11]. However, GMMs exhibit limited effectiveness in capturing the complexities of RLP distributions. An alternative approach involves using Copulas models. The study by [12] applies prominent multivariate Copulas models to simulate electric vehicle (EV) charging consumption profiles. Their findings indicate that t-Copulas outperform other functions in modeling these profiles. In recent years, the advancement of Machine Learning (ML) offers new approaches to RLP modeling. In [13, 14, 7], Generative adversarial networks (GANs) is used to either generate RLPs or PV profiles. In [15], Variational Auto-Encoder (VAE) is proposed to model the EV charging profiles. In [16], a flow-based model is proposed for RLP generation that performs better in fitting peak and probability density functions than GANs. In [17], a hybrid model VAE-GAN is proposed for the synthesis of electrical load and PV generation data. The VAE-GAN is shown to possess greater modeling capabilities compared to using a standalone GAN model. In [18], a GAN-based generative model is introduced. This model focuses on privacy instead of generating accurate profiles, transforming real-world datasets into high-quality synthetic datasets that ensure user-level privacy. A study by [19] compared the performance of a convolutional Non-linear Independent Component Estimation (NICE) model with GANs in RLP generation. The findings indicate that the convolutional NICE model produces RLPs that exhibit smaller KL divergence relative to real data, suggesting a closer approximation to the actual RLP distributions. In [20], a MultiLoad-GAN was proposed, instead of generating individual RLP, MultiLoad-GAN generates a group of synthetic RLPs which essentially capture the spatial-temporal correlations among a group of loads.
@@
Even though the above-mentioned methods show promising results, they do not include the external factors’ effect on generated RLPs (such as weather information) which is becoming more important for state-of-the-art generation methods. Conditional generation is a solution to increase the manipulability of models. In [21], cWGANs is used for probabilistic load prediction, whereas GANs is used to assist a forecaster by generating residual scenarios. In [22], a ProfileSR-GAN is proposed to upsample low-resolution RLPs to high-resolution RLPs. Furthermore, it was observed that RLPs generated with the integration of weather information lead to better generation results based on the metrics used in the research. In [23], a conditional VAE (cVAE) is proposed for representative scenario generation. In [24], two types of cVAE were employed for generating synthetic energy data. The findings from [24] suggest that the inclusion of augmented data led to a significant enhancement in building energy prediction, with accuracy improvements ranging between 12% and 18% during the experimental phase. In [25], a transferable flow-based generative model is proposed, which leverages RLP data of different households to improve the prediction of target households. In [26], a conditional MVT copula is proposed that outperforms conditional GMMs.
@@
Despite these developments, current conditional generation methods face several challenges. 1) GANs-related models are effective with discrete conditions like days and seasons but perform poorly with continuous variables such as daily or annual consumption, temperature, and irradiation [27, 28]. Moreover, models such as GANs and VAEs struggle to replicate overall statistical features because they do not directly model probability densities [29]. Copulas models handle continuous conditions well [26], but their lack of scalability makes them impractical for large datasets or high-dimensional data. 3) Flow-based models avoid the limitations mentioned above but suffer from inadequate modeling capabilities and slow convergence rates [30].
@@
In this paper, we propose a new flow-based generative model coined Full Convolutional Profile Flow (FCPFlow) [^1] You can find the code, data, and other materials of this paper at https://github.com/xiaweijie1996/Full-Convolutional-Time-Series-Flow.git, designed to address the three challenges previously discussed. FCPFlow is built upon the idea of a classical flow-based model, proposed initially in [31], designed to learn the features of RLP data efficiently. The key contributions of this paper are as follows:

### v2

@@ Abstract
Residential Load Profile (RLP) generation and prediction are critical for the operation and planning of distribution networks, particularly as diverse low-carbon technologies are increasingly integrated. This paper introduces a novel flow-based generative model, termed Full Convolutional Profile Flow (FCPFlow), which is uniquely designed for both conditional and unconditional RLP generation, and for probabilistic load forecasting. By introducing two new layers–the invertible linear layer and the invertible normalization layer–the proposed FCPFlow architecture shows three main advantages compared to traditional statistical and contemporary deep generative models: 1) it is well-suited for RLP generation under continuous conditions, such as varying weather and annual electricity consumption, 2) it shows superior scalability in different datasets compared to traditional statistical models, and 3) it also demonstrates better modeling capabilities in capturing the complex correlation of RLPs compared with deep generative models.
@@
Load Profiles, Demand Profile, Generative Model, Forecast.
@@ I Introduction
Residential Load Profiles (RLPs) have wide applications in areas such as energy supply and demand management [1], modern distribution system planning[2], and risk analysis [3]. The validity of these studies depends largely on the quality of RLPs used. However, access to RLP data is limited due to privacy [4]. RLP generation can provide effective solutions to these problems. On the one hand, distribution system operators (DSOs) rely on RLPs for refining planning decisions. Historical data inaccessibility or limitations can hinder this process, yet the generation of RLPs provides system planners with an alternative to executing more informed planning [5]. For instance, in [6], generated RLPs were used to understand consumption patterns and optimize the system planning. On the other hand, generated RLPs can function as augmented data to support high-level tasks. For example, generated RLPs or PV profiles are used to support the training of models for load prediction [7], non-intrusive load monitoring algorithms [8], and reinforcement learning [9].
@@
Traditional RLP modeling primarily employs Gaussian Mixture Models (GMMs) [10, 11]. However, GMMs exhibit limited effectiveness in capturing the complexities of RLP distributions. An alternative approach involves using Copulas models. The study by [12] applies prominent multivariate Copulas models to simulate electric vehicle (EV) charging consumption profiles. Their findings indicate that t-Copulas outperform other functions in modeling these profiles. In recent years, the advancement of Machine Learning (ML) offers new approaches to RLP modeling. In [13, 14, 7], Generative adversarial networks (GANs) are used to either generate RLPs or PV profiles. In [15], Variational Auto-Encoder (VAE) is proposed to model EV charging profiles. In [16], a flow-based model is proposed for RLP generation that performs better in fitting peak and probability density functions than GANs. In [17], a hybrid model VAE-GAN is proposed for the synthesis of electrical load and PV generation data. The VAE-GAN is shown to possess greater modeling capabilities compared to using a standalone GAN model. In [18], a GAN-based generative model is introduced. This model focuses on privacy instead of generating accurate profiles, transforming real-world datasets into high-quality synthetic datasets that ensure user-level privacy. A study by [19] compared the performance of a convolutional Non-linear Independent Component Estimation (NICE) model with GANs in RLP generation. The findings indicate that the convolutional NICE model produces RLPs that exhibit smaller KL divergence relative to real data, suggesting a closer approximation to the actual RLP distributions. In [20], a MultiLoad-GAN was proposed, instead of generating individual RLP, MultiLoad-GAN generates a group of synthetic RLPs which essentially capture the spatial-temporal correlations among a group of loads.
@@
Even though the above-mentioned methods show promising results, they do not include the effects of external factors on generated RLPs (such as weather information), which is becoming more important for state-of-the-art generation methods. Conditional generation is a solution to increase the manipulability of models. In [21], cWGANs are used for probabilistic load prediction, whereas GANs are used to assist a forecaster by generating residual scenarios. In [22], a ProfileSR-GAN is proposed to upsample low-resolution RLPs to high-resolution RLPs. Furthermore, it was observed that RLPs generated with the integration of weather information lead to better generation results based on the metrics used in the research. In [23], a conditional VAE (cVAE) is proposed for representative scenario generation. In [24], two types of cVAE were employed for generating synthetic energy data. The findings from [24] suggest that including augmented data significantly enhanced building energy prediction, with accuracy improvements ranging between 12% and 18% during the experimental phase. In [25], a transferable flow-based generative model is proposed, which leverages RLP data of different households to improve the prediction of target households. In [26], a conditional MVT copula is proposed that outperforms conditional GMMs.
@@
Despite these developments, current conditional generation methods face several challenges. 1) GANs-related models are effective with discrete conditions like days and seasons but perform poorly with continuous variables such as daily or annual consumption, temperature, and irradiation [27, 28]. Moreover, models such as GANs and VAEs struggle to replicate overall statistical features because they do not directly model probability densities [29]. Copulas models handle continuous conditions well [26], but their lack of scalability makes them impractical for large datasets or high-dimensional data. 3) Flow-based models avoid the above-mentioned limitations but suffer from inadequate modeling capabilities and slow convergence rates [30].
@@
In this paper, we propose a new flow-based generative model architecture coined Full Convolutional Profile Flow (FCPFlow) [^1] You can find the code, data, and other materials of this paper at https://github.com/xiaweijie1996/Full-Convolutional-Time-Series-Flow.git, designed to address the three challenges previously discussed. The proposed FCPFlow architecture is built upon the idea of a classical flow-based model, proposed initially in [31], but designed to learn the features of RLP data efficiently. The key contributions of this paper are as follows: